The Perimeter of a playing field is 504 yards and the length is 2 times the width minus six. What is the length and width?
Answer by palanisamy(496)

Let the width be x.
Given,the length is 2 times the width minus six.
So the length = 2x-6
Perimeter is 2(x+2x-6) = 504
2(3x-6) = 504
6x-12 = 504
6x = 504+12
6x = 516
x = 516/6
x = 86
Therefore width = 86 yards
length= 2x-6 = 2*86-6
= 172-6
= 166 yards